# Suppressing Correlated Noise in Quantum Computers via Context-Aware Compiling

### v1

@@ I Introduction
Among the variety of possible errors, coherent quantum noise can harm computation quadratically more than incoherent (stochastic) errors. Therefore, their accurate characterization and suppression has become an important goal [9, 10]. Worse, some of these errors exhibit correlations across several qubits, thus spreading errors faster and making them particularly challenging to handle in error correction protocols [11, 12]. In this paper, we present methods and experimental results for suppressing a wide range of correlated coherent errors, using the context of the hardware and the circuits that run on it. Throughout, we present a comprehensive characterization of several sources of errors, and discuss different strategies that suit each one.
In the near-term, the intersection of error suppression and error mitigation deserves a special attention. Error mitigation techniques typically use additional samples of the circuit and post-processing to improve the accuracy of the results. In leading protocols such as Probabilistic Error Cancellation (PEC) [4, 5] and Probabilistic Error Amplification (PEA) [13], circuits are arranged in layers of gates that are subsequently twirled [14, 15, 16, 17, 18] by inserting random Pauli gates between them without changing the overall logic of the circuit. This twirling simplifies the structure of the noise in circuit layers and makes it amenable to error mitigation. The cost of error mitigation with PEC, as determined by its sampling overhead, grows exponentially with circuit size, and the base of the exponential is determined by the overall noise in the circuit. Therefore, even modest improvements in the errors can have significant impact on the total runtime of an error mitigated computation.
@@ II.2 Dynamical decoupling
While the above example illustrated a simple single-qubit DD sequence for suppressing single-qubit $Z$ errors, finding the sequence that effectively decouples a multi-qubit system with an arbitrary Hamiltonian to an arbitrary order in $\tau$ can be challenging [21]. In particular, to suppress a coherent two-qubit $ZZ$ error in addition to coherent single qubit $Z$ errors—i.e. those generated by $H_{11}$ (1)—it is necessary to apply DD pulses on two qubits in a staggered fashion (see Fig. 1). This is because the $X\otimes X$ from DD commutes with the unwanted $Z\otimes Z$, and there will not be any cancellation. However, staggering the pulses ensures that there are periods of evolution accumulating opposite phases that eventually cancel and fully suppress the unwanted interactions.
While the implementation of DD in memory-like experiments, i.e., where the objective is to preserve a quantum state over an idling time, has been successful [41], its integration into arbitrary quantum circuits requires meticulous attention to the context. Specifically, the presence of neighboring single or two-qubit gates necessitates careful alignments that maximize error suppression over all qubits. This becomes increasingly challenging as the complexity of the circuit scales. We formulate this problem as a constrained graph coloring one, and show that depending on crosstalk terms of the device and the presence of gate spectators at any given time, we may need $3$ or more colors even when the qubit graph is bipartite.
@@ III.1 Layered circuits and Pauli twirling
Here, we consider the problem of performing error suppression at scale. Specifically, we consider error suppression as a stage prior to error mitigation to reduce the mitigation overhead. In a typical workflow of error mitigation protocols such as PEC and PEA, an arbitrary circuit is stratified into alternating layers of single-qubit and two-qubit gates (see Fig. 2). Afterwards, the two-qubit gates are twirled, that is random single qubit gates are applied before and after the gates without altering the logical operation of the circuit. Twirling then simplifies the characterization and mitigation of errors [14, 15, 16, 17, 18]. Here, we focus on the specific case of Pauli twirling of Clifford two-qubit gate layers and assume that errors on single-qubit gates are gate-independent. Note that as any unitary operation can be realized by a gateset consisting of an entangling two-qubit gate and arbitrary single qubit rotations, any circuit can be stratified into alternating layers of arbitrary single qubit gates and Clifford two-qubit gates [45, 46].
@@ III.2 Context-dependent errors
The first example we consider (case I in Fig. 3) is when there are two adjacent idle qubits for some time $\tau$. In this case, the qubits are affected by a coherent two-qubit error Hamiltonian $H_{11}$ (1). Therefore, the total error is given by $U_{11}=R_{zz}(\theta)\cdot[R_{z}(-\theta)\otimes R_{z}(-\theta)]$ (2), where $\theta=\nu\tau$. When these error are followed by a Pauli twirling layer and an arbitrary two-qubit gate, the $R_{z}$ terms can be compensated in the single qubit layer, and the $R_{zz}$ term can be moved past the twirling gates and compensated by the two-qubit gate that follows. Alternatively, context-aware staggered DD sequences can also suppress these errors. However, if DD is applied on individual qubits without considering the context, the $X$ gates in DD align and cannot fully suppress the $U_{11}$ error. Specifically, while aligned DD cancels the $R_{z}(-\theta)$ terms, it cannot suppress $R_{zz}(\theta)$. In this case, one can again compensate the remaining error term in the following two-qubit gates. The results for all these suppression techniques for a varying idling time of $d\tau$ are experimentally demonstrated in Fig. 3c. The remarkable improvement observed by error compensation validates our model for the major source of coherent noise in this experiment. The difference in the asymptotic decay rates of the error compensation and staggered DD curves is indicative of the presence of temporally correlated incoherent noise, which DD is effective at suppressing.
@@ 5       $S^{\prime}\leftarrow$ ApplyDDSeqByColor($S$, $C$, $L_{\text{DD}}$);
6
7Function CollectJointDelays($S$, $G$, $D_{\text{min}}$):
@@ 16                         $w$leftarrow delay_interval ;
23
24Function ColorGraph($I$, $G$, $S$, $\mathcal{H}$):
@@ IV.2 Context-aware error compensation (CA-EC)
1 $compensation\_1q\leftarrow\{q:\emptyset\textnormal{ for n in }C.nodes()\}$;
$compensation\_2q\leftarrow\{e:\emptyset\textnormal{ for e in }C.edges()\}$; /* Split circuit into layers with either only 1q or only 2q gates */
@@ VI Related work
Dynamical decoupling has a long history in quantum systems [19], though the vast majority of demonstrations have been in small systems or via manual pulse insertion [65, 24, 66]. Recently, several works have applied DD with success to larger multi-qubit circuits, but in a uniform fashion across all qubits [23, 67], which fails to address non-local errors. More recently, several experiments have demonstrated crosstalk suppression using staggered dynamical decoupling [68, 41], building on theoretical proposals from decades ago [55, 54]. However these have been manual with no compiler support. The idea of adapting DD sequences to different parts of the circuit has gained traction recently [69, 70, 71] but these often require tuning the placements via many additional circuit executions and are not informed by the physical properties of the noise. Suppressing crosstalk by changing the circuit schedule or tuning qubit frequencies have also been studied in Ref. [72, 73], though with limited experimental evaluation and non-scalable compilation via exact solvers. Lastly, there has been some research on adapting circuits to combat coherent errors, though mainly focused on gate errors and not correlated spectator errors. Ref. [74] alternates the physical implementation of a gate within a circuit to get an overall suppression. Similarly, Ref. [75] re-compiles the circuit after considering errors arising from gates on a flux-tunable architecture, and evaluate the proposal numerically.

### v2

@@ I Introduction
Among the variety of possible errors, coherent quantum noise can harm computation quadratically more than incoherent (stochastic) errors. Therefore, their accurate characterization and suppression has become an important goal [9, 10, 11]. Worse, some of these errors exhibit correlations across several qubits, thus spreading errors faster and making them particularly challenging to handle in error correction protocols [12, 13]. In this paper, we present methods and experimental results for suppressing a wide range of correlated coherent errors, using the context of the hardware and the circuits that run on it. Throughout, we present a comprehensive characterization of several sources of errors, and discuss different strategies that suit each one.
In the near-term, the intersection of error suppression and error mitigation deserves a special attention. Error mitigation techniques typically use additional samples of the circuit and post-processing to improve the accuracy of the results. In leading protocols such as Probabilistic Error Cancellation (PEC) [4, 5] and Probabilistic Error Amplification (PEA) [14], circuits are arranged in layers of gates that are subsequently twirled [15, 16, 17, 18, 19, 20] by inserting random Pauli gates between them without changing the overall logic of the circuit. This twirling simplifies the structure of the noise in circuit layers and makes it amenable to error mitigation. The cost of error mitigation with PEC, as determined by its sampling overhead, grows exponentially with circuit size, and the base of the exponential is determined by the overall noise in the circuit. Therefore, even modest improvements in the errors can have significant impact on the total runtime of an error mitigated computation.
@@ II.2 Dynamical decoupling
While the above example illustrated a simple single-qubit DD sequence for suppressing single-qubit $Z$ errors, finding the sequence that effectively decouples a multi-qubit system with an arbitrary Hamiltonian to an arbitrary order in $\tau$ can be challenging [23]. In particular, to suppress a coherent two-qubit $ZZ$ error in addition to coherent single qubit $Z$ errors—i.e. those generated by $H_{11}$ (1)—it is necessary to apply DD pulses on two qubits in a staggered fashion (see Fig. 1). This is because the $X\otimes X$ from DD commutes with the unwanted $Z\otimes Z$, and there will not be any cancellation. However, staggering the pulses ensures that there are periods of evolution accumulating opposite phases that eventually cancel and fully suppress the unwanted interactions [43, 44].
While the implementation of DD in memory-like experiments, i.e., where the objective is to preserve a quantum state over an idling time, has been successful [43, 44], its integration into arbitrary quantum circuits requires meticulous attention to the context. Specifically, the presence of neighboring single or two-qubit gates necessitates careful alignments that maximize error suppression over all qubits. This becomes increasingly challenging as the complexity of the circuit scales. We formulate this problem as a constrained graph coloring one, and show that depending on crosstalk terms of the device and the presence of gate spectators at any given time, we may need $3$ or more colors even when the qubit graph is bipartite.
@@ III.1 Layered circuits and Pauli twirling
Here, we consider the problem of performing error suppression at scale. Specifically, we consider error suppression as a stage prior to error mitigation to reduce the mitigation overhead. In a typical workflow of error mitigation protocols such as PEC and PEA, an arbitrary circuit is stratified into alternating layers of single-qubit and two-qubit gates (see Fig. 2). Afterwards, the two-qubit gates are twirled, that is random single qubit gates are applied before and after the gates without altering the logical operation of the circuit. Twirling then simplifies the characterization and mitigation of errors [15, 16, 17, 18, 19, 20]. Here, we focus on the specific case of Pauli twirling of Clifford two-qubit gate layers and assume that errors on single-qubit gates are gate-independent. Note that as any unitary operation can be realized by a gateset consisting of an entangling two-qubit gate and arbitrary single qubit rotations, any circuit can be stratified into alternating layers of arbitrary single qubit gates and Clifford two-qubit gates [48, 49].
@@ III.2 Context-dependent errors
The first example we consider (case I in Fig. 3) is when there are two adjacent idle qubits for some time $\tau$. In this case, the qubits are affected by a coherent two-qubit error Hamiltonian $H_{11}$ (1). Therefore, the total error is given by $U_{11}=R_{zz}(\theta)\cdot[R_{z}(-\theta)\otimes R_{z}(-\theta)]$ (2), where $\theta=\nu\tau$. When these error are followed by a Pauli twirling layer and an arbitrary two-qubit gate, the $R_{z}$ terms can be compensated in the single qubit layer, and the $R_{zz}$ term can be moved past the twirling gates and compensated by the two-qubit gate that follows. Alternatively, context-aware staggered DD sequences can also suppress these errors [43, 44]. However, if DD is applied on individual qubits without considering the context, the $X$ gates in DD align and cannot fully suppress the $U_{11}$ error. Specifically, while aligned DD cancels the $R_{z}(-\theta)$ terms, it cannot suppress $R_{zz}(\theta)$. In this case, one can again compensate the remaining error term in the following two-qubit gates. The results for all these suppression techniques for a varying idling time of $d\tau$ are experimentally demonstrated in Fig. 3c. The remarkable improvement observed by error compensation validates our model for the major source of coherent noise in this experiment. The difference in the asymptotic decay rates of the error compensation and staggered DD curves is indicative of the presence of temporally correlated incoherent noise, which DD is effective at suppressing.
@@ 11       $S^{\prime}\leftarrow$ ApplyDDSeqByColor($S$, $C$, $L_{\text{DD}}$);
12
13Function CollectJointDelays($S$, $G$, $D_{\text{min}}$):
@@ 22                         $w$leftarrow delay_interval ;
29
30Function ColorGraph($I$, $G$, $S$, $\mathcal{H}$):
@@ IV.2 Context-aware error compensation (CA-EC)
2
3
[…]
5 $compensation\_2q\leftarrow\{e:\emptyset\textnormal{ for e in }C.edges()\}$;
/* Split circuit into layers with either only 1q or only 2q gates */
@@ VI Related work
Dynamical decoupling has a long history in quantum systems [21], though the vast majority of demonstrations have been in small systems or via manual pulse insertion [67, 26, 68]. Recently, several works have applied DD with success to larger multi-qubit circuits, but in a uniform fashion across all qubits [25, 69], which fails to address non-local errors. More recently, several experiments have demonstrated crosstalk suppression using staggered dynamical decoupling [70, 43], building on theoretical proposals from decades ago [57, 56]. However these have been manual with no compiler support. The idea of adapting DD sequences to different parts of the circuit has gained traction recently [71, 72, 73, 74] but these often require tuning the placements via many additional circuit executions and are not informed by the physical properties of the noise. Suppressing crosstalk by changing the circuit schedule or tuning qubit frequencies have also been studied in Ref. [75, 76], though with limited experimental evaluation and non-scalable compilation via exact solvers. Lastly, there has been some research on adapting circuits to combat coherent errors, though mainly focused on gate errors and not correlated spectator errors [77]. Ref. [78] alternates the physical implementation of a gate within a circuit to get an overall suppression. Similarly, Ref. [79] re-compiles the circuit after considering errors arising from gates on a flux-tunable architecture, and evaluate the proposal numerically.